# MA4DIV: Multi-Agent Reinforcement Learning for Search Result Diversification

### v2

@@ 4. The MA4DIV Model
In this section, we first describe how to formulate the search results diversification problem as a Dec-POMDP by introducing the basic elements in the MA4DIV framework. Then, we present the architecture of the MA4DIV framework, as shown in Figure 1, and elaborate the implementation of the Agent Network, Ranking Process, and Mixing Network. Finally, we will give an algorithm to train the MA4DIV.
@@ 4.1. Essential Elements of MA4DIV
In Section 3.1, we define the process of diversity ranking as a Dec-POMDP. Next, we introduce the essential elements of Dec-POMDP specified to our MA4DIV:
@@
Observation $\mathcal{O}$: Observation represents the localized information that each agent receives. Each agent makes decisions based on its respective observation $\bm{o_{i}}$. The defination of $\bm{o_{i}}$ is shown as Equation (6).
(6) $\bm{o_{i}}=\left\{\mathbf{q},\mathbf{d_{i}},\textbf{D}\right\}$
where $\mathbf{q}$ is embedding vector of the given query, $\mathbf{d_{i}}$ is embedding vector of document $i$ and $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$ represents the embedding vectors of all documents.
@@
Actions $\mathcal{A}$: The action space $a_{i}\in\mathcal{A}$ of the MA4DIV algorithm is defined as a set of discrete ranking scores $s_{i}\in\{1,\dots,\mathcal{|A|}\}$. When making decisions, agent $i$ will select an action $a_{i}$ from $\mathcal{A}$. The corresponding score $s_{i}$ serves as the basis for agent $i$ (document $i$) to assess the level of diversity to itself.
@@
Reward $\mathcal{R}$: The reward function should guide the parameter updates of the model to achieve diversity of ranked documents. In search result diversification, there are some evaluation metrics, such as $\alpha$-NDCG, are used to evaluate the diversity of a ranked document list. So it’s natural to consider these metrics as reward functions. In this paper, we define the reward function as $\alpha$-NDCG at top $k$ positions:
@@ 4.2. Agent Network
In Figure 1, the green modules represent the Agent Network, which share parameters with each other. On the left side of Figure 1 is the detailed structure of the Agent Network. It can be seen that we concatenate $\{\mathbf{q},\mathbf{d_{i}},\mathbf{e_{i}}\}$, which is precisely the observation for agent $i$ defined in Equation (6). And $\mathbf{e_{i}}$ is a vector with high-level cross feature between document $i$ and all other documents, computed by Multi-Head Self-Attention (MHSA) module:
MHSA further enhances the model capacity by applying the above process multiple times in parallel, which is called multi-head mechanism. Suppose there are $h$ heads, the output after multi-head is $\mathbf{O}=[\mathbf{O}_{1};\mathbf{O}_{2};\cdots;\mathbf{O}_{h}]\mathbf{W}_{O}$, where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation, and $\mathbf{O}_{i}$ and $\mathbf{W}_{O}$ are the output of the $i$-th head and a learnable weight matrix with dimensions $h\cdot d_{v}\times d$, respectively. Finally, $\mathbf{O}$ is reshaped to output $\mathbf{e_{i}}$ as Equation (11):
(11) $\{\mathbf{e_{1}},\dots,\mathbf{e_{i}},\dots,\mathbf{e_{n}}\}=\mathbf{MHSA}([\mathbf{d_{1}},\dots,\mathbf{d_{n}}])$
According to (Pang et al., 2020), the ranking model should satisfy the permutation invariance property which is described as following Defination:
Definition 0 (Permutation Invariance).
A method to construct a ranking model in accordance with Permutation Invariance involves initially attributing scores to the documents via a permutation equivariant scoring function, followed by ordering based on these scores. Pang et al. (2020) also proved that the multi-head attention block is permutation equivariant. Therefore, the MHSA module is used to compute higher-level cross feature vectors between documents.
@@
(12) $\{Q_{i}^{a_{1}},\dots,Q_{i}^{a_{m}},\dots,Q_{i}^{a_{n}}\}$ $=\mathbf{MLP}\left(\{\mathbf{q},\mathbf{d_{i}},\mathbf{e_{i}}\}\right)$
@@
Specifically, with a probability of $1-\varepsilon$, the agent $i$ chooses the action-value $Q_{i}^{a_{*}}$ which has the highest estimated value ($\max Q_{i}^{a_{m}}$) for the current observation $\bm{o_{i}}$, and the corresponding chosen action is $a_{*}^{i}=\arg\max_{a_{m}}Q_{i}^{a_{m}},\forall m\in\{1,\dots,|\mathcal{A}|\}$. This is the exploitation part, where the agent makes the best decision based on the current model. On the other hand, with a probability of $\varepsilon$, the agent $i$ chooses an action-value uniformly at random from $\mathcal{|A|}$ action-values. This is the exploration part, where the agents try to discover new permutations of the candidate documents set. Mathematically, this process can be written as Equation (14):
@@ 4.3. Ranking Process
In the ranking process, the agents receive a query q and the associated documents $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$. As shown in Figure 1, each agent $i$ select its own action-value $Q_{i}^{a_{*}}$ and the corresponding action $a_{*}^{i}$ according to Equation (14). Because the action space $\mathcal{A}$ is defined as discrete ranking scores, that is, each action $a_{m}$ corresponds to a ranking score. So the ranking scores of all documents can be obtained as $\mathbf{scores}=\{s_{1},\dots,s_{n}\}$. Finally, a ranked document list $\mathbf{D_{ranked}}$ is obtained based on the descending order of $\mathbf{scores}$.
@@
2 Select Actions: Each agent $i$ select action-value $Q_{i}^{a_{*}}$ to get $\{Q_{1}^{a_{*}},\dots,Q_{i}^{a_{*}},\dots,Q_{n}^{a_{*}}\}$ and $\{a_{*}^{1},\dots,a_{*}^{i},\dots,a_{*}^{n}\}$ base on Equation (14).
3 Getting Scores: Get the corresponding ranking scores $\mathbf{scores}=\{s_{1},\dots,s_{i},\dots,s_{n}\}$.
@@ 4.4. Value-Decomposition for MA4DIV
The MARL algorithms based on value decomposition, such as (Sunehag et al., 2017; Rashid et al., 2018; Wang et al., 2020b), consider that there is the global utility function $Q_{tot}^{*}$ can be decomposed into the action-value functions $Q_{i}^{a_{*}}$ of all individual agents. $Q_{tot}^{*}$ represents the action-value utility of the joint-action $\mathbf{a}$ under the joint-observation information $\mathbf{o}$, and is also a scalar value. The monotonic decomposition relation can be expressed as Equation (15):
@@ 4.5. Mixing & Hyper Network Structure
To convert the individual action-value functions $Q_{i}^{a_{*}}$ into the global utility function $Q_{tot}^{*}$ and ensure compliance with Equation (16), an architecture consisting of a mixing network and a set of hypernetworks is designed. The right part of Figure 1 illustrates the detail structure.
@@
Hypernetworks, several fully connected networks, play a role in generating mixing network weights. Firstly, each hypernetwork takes global state $\bm{s}$ as input and generate the vectors of $\mathbf{W1}$, $\mathbf{B1}$, $\mathbf{W2}$, B2 respectively. Then, the vectors are reshaped into matrices of appropriate size as parameters for the mixing network. Lastly, to achieve the monotonicity constraint of Equation (16), $\mathbf{W1}$, $\mathbf{W2}$ are limited to non-negative by an absolute activation function. However, biases $\mathbf{B1}$, B2 do not need to be set to non-negative because biases has nothing to do with the partial derivatives between $Q_{tot}^{*}$ and $Q_{i}^{a_{*}}$. The hypernetworks operate as follows:
(18) $\mathbf{W1}&=\mathbf{Reshape}\left(\left|\mathbf{FC_{{\scriptscriptstyle W1}}}\left(\bm{s}\right)\right|\right),\mathbf{B1}=\mathbf{Reshape}\left(\mathbf{FC_{{\scriptscriptstyle B1}}}\left(\bm{s}\right)\right)\\ \mathbf{W2}&=\mathbf{Reshape}\left(\left|\mathbf{FC_{{\scriptscriptstyle W2}}}\left(\bm{s}\right)\right|\right),\text{B2}=\mathbf{Reshape}\left(\mathbf{FC_{{\scriptscriptstyle B2}}}\left(\bm{s}\right)\right)\end{aligned}\end{gathered}$
@@ 4.6. Training Process of MA4DIV
In this section, we introduce the complete training process for MA4DIV. The parameters of models are $\Theta=\{\theta,\psi,\phi\}$ where $\theta$ represents the parameters of agent network, $\psi$ and $\phi$ are the parameters of mixing network and hypernetworks respectively. The model parameters $\Theta$ is updated on $N$ training data with the form of $\{(q^{k},D^{k}),J^{k}\}_{k=1}^{N}$.
Algorithm 2 shows the training pseudo-code. At the beginning, parameters $\Theta$ are initialized randomly and the replay buffer $\mathcal{M}$ is set to $\varnothing$. A whole training epoch consists of two parts. The first part is the stage of data generation. Given data $\mathbf{q}=q^{i}$, $\mathbf{D}=D^{i}$, with subtopic labels $J^{i}$, tuple $\mathcal{T}$ in the episode is obtained based on the definition in Equation (6, 7, 8). Then, store the tuple $\mathcal{T}$ in replay buffer. Each epoch can generate a kind of permutation of all documents $\{D^{k}\}_{k=1}^{N}$ in all training data. The second part is updating models with parameters $\Theta$ for $N\_update$ times. And the update is performed by sampling data through a mini-batch approach from $\mathcal{M}$. The whole network is trained end-to-end by minimizing the loss shown in Equation (19) and (20) to maximize $Q_{tot}$.
(20) $y^{tot}$ $=\mathcal{R}+\gamma\max_{a^{\prime}}Q_{tot}(\textbf{o}^{\prime},\textbf{a}^{\prime},\bm{s}^{\prime};\Theta^{\prime})$
$=\mathcal{R}$
@@
9             Collect a tuple $\mathcal{T}$ = ($\mathbf{o},\bm{s},\mathbf{a},\mathcal{R},\mathbf{o^{\prime}},\bm{s^{\prime}}$) during this episode.
@@
Table 1. Comparisons of datasets TREC and DU-DIV.
Dataset Source No. of Queries No. of Documents per Query No. of Subtopics Representation
TREC TREC 2009$\sim$2012 Web Track 198 average 211 docs/query maximum 7 subtopics/doc Doc2vec (Le and Mikolov, 2014)
DU-DIV Baidu 4473 15 docs/query total 50 subtopics BERT (Devlin et al., 2018)
@@
Table 2. Hyperparemeters of MA4DIV on different dataset.
Dataset $|\mathcal{A}|$ $L$ $H$ $z$
TREC 30 1 4 64
DU-DIV 15 1 4 256
@@ 5. Experiments
(RQ.2) How does MA4DIV perform on a large-scale industrial dataset comparing to existing baselines?
(RQ.3) How efficient is MA4DIV in training process?
5.1. Datasets
We conduct experiments to address above research questions on two datasets, TREC 2009$\sim$2012 Web Track datasets and DU-DIV dataset. TREC 2009$\sim$2012 Web Track datasets are publicly available and Xia et al. (2017) first uses them for training and evaluating diversified search models. Almost all the following works, such as $\mathrm{M^{2}Div}$ (Feng et al., 2018), DALETOR (Yan et al., 2021), MO4SRD (Yu, 2022), etc., conduct experiments on these datasets. As a widely used test collection for diversified search, the TREC Web track dataset only contains 198 queries, and previous studies often use Doc2vec (Le and Mikolov, 2014) to obtain the document embeddings on this dataset. In this work, to ensure a fair and meaningful comparison with existing work, we first compare MA4DIV with existing baselines on the this relatively small dataset. Then, we further conduct extensive experiments on a new and larger scale diversity dataset called DU-DIV, which is constructed based on real user data in Baidu search engine.
The details of these two datasets are shown in Table 1. The DU-DIV dataset has 4473 queries, far more than the TREC datasets. The documents in the DU-DIV dataset are the top-15 documents retrieved by the Baidu search engine for each query, which makes this dataset particularly useful for the diversified ranking task in which we need to rerank these top documents for diversity. In the TREC web track datasets, each document contains a maximum of 7 subtopics. In the DU-DIV dataset, we predefined 50 subtopics, such as law, healthcare, education, etc. Each document corresponds to one or more of the predefined subtopics. Moreover, while previous studies on the TREC web track dataset often use the Doc2vec model to represent queries and documents in vector with $L\textrm{=}100$ dimensions, the DU-DIV dataset utilizes the more powerful model BERT (Devlin et al., 2018) to encode queries and documents into vectors with $L\textrm{=}1024$ dimensions.
5.2. Experimental Settings
Considering that TREC datasets contain only 198 queries, this is prone to overfitting for existing deep learning models. In addition, the average number of documents related to each query is about 211, but a large number of these documents do not contain any subtopics. Therefore, we reshape TREC dataset. That is, from each of the original document lists in TREC (198 lists in total, with an average of 211 documents per query), we subsample 30 documents at a time to form a new document list. This process runs multiple times over a document list associated with the same query, resulting in a reshaped dataset with a total of 6,232 document lists (198 unique queries). And it’s crucial to ensure that an appropriate number of the 30 documents sampled contain at least one subtopic.

### v3

@@ 4. The MA4DIV Model
In this section, we first describe how to formulate the search results diversification problem as a Co-MARL by introducing the basic elements in the MA4DIV framework. Then, we present the architecture of the MA4DIV framework, as shown in Figure 1, and elaborate the implementation of the Agent Network, Ranking Process, and Mixing Network. Finally, we will give an algorithm to train the MA4DIV.
@@ 4.1. Essential Elements of MA4DIV
In Section 3.1, we define the process of diversity ranking as a Co-MARL. Next, we introduce the essential elements of Co-MARL specified to our MA4DIV:
@@
Observation $\mathcal{O}$: Observation represents the information that each agent receives. Each agent makes decisions based on its respective observation $\bm{o_{i}}$. The definition of $\bm{o_{i}}$ is shown as Equation (2).
(2) $\bm{o_{i}}=\left\{\mathbf{q},\textbf{D},\mathbf{d_{i}}\right\}$
where $\mathbf{q}$ is embedding vector of the given query, $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$ represents the embedding vectors of all documents and $\mathbf{d_{i}}$ is embedding vector of document $i$ to make $\bm{o_{i}}$ be personalized to each agent.
Actions $\mathcal{A}$: The action space $a_{i}\in\mathcal{A}$ of each agent is corresponding to a set of integer ranking scores $s_{i}\in\{1,\dots,\mathcal{|A|}\}$, $\mathcal{|A|}$ is the dimension of the action space. When making decisions, agent $i$ will select an action $a_{i}$ from $\mathcal{A}$. And the action $a_{i}$ corresponds to a ranking score $s_{i}$, which will serve as the basis for agent $i$ (document $i$) to assess the level of diversity to itself, and thus, be used in generating the final ranking list.
@@
Reward $\mathcal{R}$: The reward function should guide the parameter updates of the model to achieve diversity of ranked documents. In search result diversification, there are some evaluation metrics, such as $\alpha$-NDCG, are used to evaluate the diversity of a ranked document list. So it’s natural to consider these metrics as reward functions. In this paper, we define the reward function as $\alpha$-NDCG@k:
@@ 4.2. Agent Network
In Figure 1, the green modules represent the Agent Network for an agent $i$, which share parameters with each other. On the left side of Figure 1 is the detailed structure of the Agent Network. It can be seen that we concatenate $\{\mathbf{q},\mathbf{d_{i}},\mathbf{e_{i}}\}$, which is precisely obtained from the observation for agent $i$ defined in Equation (2). And $\mathbf{e_{i}}$ is a vector with high-level cross feature between document $i$ and all other documents, computed by Multi-Head Self-Attention (MHSA) module. The detail of MHSA and the reason why we use MHSA can be found in Appendix C. Finally, the features of all documents $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$ is reshaped to output $\mathbf{e_{i}}$ as Equation (6):
(6) $\{\mathbf{e_{1}},\dots,\mathbf{e_{i}},\dots,\mathbf{e_{n}}\}=\mathbf{MHSA}(\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\})$
@@
(7) $\{Q_{i}^{a_{1}},\dots,Q_{i}^{a_{m}},\dots,Q_{i}^{a_{|\mathcal{A}|}}\}$ $=\mathbf{MLP}\left(\{\mathbf{q},\mathbf{d_{i}},\mathbf{e_{i}}\}\right)$
@@
Specifically, with a probability of $1-\varepsilon$, the agent $i$ chooses the action-value $Q_{i}^{a_{*}}$ which has the highest estimated value ($\max Q_{i}^{a_{m}}$) for the current observation $\bm{o_{i}}$, and the corresponding chosen action is $a_{i}^{*}=\arg\max_{a_{m}}Q_{i}^{a_{m}},\forall m\in\{1,\dots,|\mathcal{A}|\}$. This is the exploitation part, where the agent makes the best decision based on the current model. On the other hand, with a probability of $\varepsilon$, the agent $i$ chooses an action-value uniformly at random from $\mathcal{|A|}$ action-values. This is the exploration part, where the agents try to discover new permutations of the candidate documents set. Mathematically, this process can be written as Equation (9):
@@ 4.3. Ranking Process
In the ranking process, the agents receive a query q and the associated documents $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$. As shown in Figure 1, each agent $i$ select its own action-value $Q_{i}^{a_{*}}$ and the corresponding action $a_{i}^{*}$ according to Equation (9). Because the action space $\mathcal{A}$ is defined as integer ranking scores, that is, each action $a_{m}$ corresponds to a ranking score. So the ranking scores of all documents can be obtained as $\mathbf{scores}=\{s_{1},\dots,s_{n}\}$. Finally, a ranked document list $\mathbf{D_{ranked}}$ is obtained based on the descending order of $\mathbf{scores}$.
@@
2 Select Actions: Each agent $i$ select action-value $Q_{i}^{a_{*}}$ to get $\{Q_{1}^{a_{*}},\dots,Q_{i}^{a_{*}},\dots,Q_{n}^{a_{*}}\}$ based on Equation (9), and the corresponding joint-action $\mathbf{a}^{*}=\{a_{1}^{*},\dots,a_{i}^{*},\dots,a_{n}^{*}\}$.
3 Getting Scores: Obtain the $\mathbf{scores}=\{s_{1},\dots,s_{i},\dots,s_{n}\}$ corresponding to the joint-action $\mathbf{a}^{*}$.
@@ 4.4. Value-Decomposition for MA4DIV
The MARL algorithms based on value decomposition, such as (Sunehag et al., 2017; Rashid et al., 2018; Wang et al., 2020b), consider that there is the global utility function $Q_{tot}^{*}$ can be decomposed into the action-value functions $Q_{i}^{a_{*}}$ of all individual agents. $Q_{tot}^{*}$ represents the action-value utility of the joint-action $\mathbf{a}^{*}=\{a_{1}^{*},\dots,a_{n}^{*}\}$ under the joint-observation information $\mathbf{o}=\{\bm{o_{1}},\dots,\bm{o_{n}}\}$, and is also a scalar value. The monotonic decomposition relation can be expressed as Equation (10):
@@ 4.5. Mixing & Hyper Network Structure
To convert the individual action-value functions $Q_{i}^{a_{*}}$ into the global utility function $Q_{tot}^{*}$ and ensure compliance with Equation (11), an architecture consisting of a mixing network and a set of hypernetworks are designed. The right part of Figure 1 illustrates the detail structure.
@@
(12) $Q_{tot}^{*}=\mathbf{W_{2}}\cdot\mathbf{Elu}\left(\mathbf{W_{1}}\cdot[Q_{1}^{a_{*}},\dots,Q_{n}^{a_{*}}]+\mathbf{B_{1}}\right)+B_{2}$
Hypernetworks, several fully connected networks, play a role in generating mixing network weights. Firstly, each hypernetwork takes global state $\bm{s}$ as input and generate the vectors of $\mathbf{W_{1}}$, $\mathbf{B_{1}}$, $\mathbf{W_{2}}$, $B_{2}$ respectively. Then, the vectors are reshaped into matrices of appropriate size as parameters for the mixing network. Lastly, to achieve the monotonicity constraint of Equation (11), $\mathbf{W_{1}}$, $\mathbf{W_{2}}$ are limited to non-negative by an absolute activation function. However, biases $\mathbf{B_{1}}$, $B_{2}$ do not need to be set to non-negative because biases has nothing to do with the partial derivatives between $Q_{tot}^{*}$ and $Q_{i}^{a_{*}}$. The hypernetworks operate as follows:
@@ 4.6. Training Process of MA4DIV
In this section, we introduce the complete training process for MA4DIV. The parameters of models are $\Theta=\{\theta,\psi,\phi\}$ where $\theta$ represents the parameters of agent network, $\psi$ and $\phi$ are the parameters of mixing network and hypernetworks respectively. The model parameters $\Theta$ is updated on $N$ training data with the form of $\{(q^{k},D^{k}),J^{k}\}_{k=1}^{N}$ mentioned in Section 3.2.
@@
9             Collect a tuple $\mathcal{T}$ = ($\mathbf{o},\bm{s},\mathbf{a},\mathcal{R}$) during this episode.
@@
Algorithm 2 shows the training pseudo-code. At the beginning, parameters $\Theta$ are initialized randomly and the replay buffer $\mathcal{M}$ is set to $\varnothing$. A whole training epoch consists of two parts. The first part is the stage of data generation. Given data $\mathbf{q}=q^{i}$, $\mathbf{D}=D^{i}$, with subtopic labels $J^{i}$, tuple $\mathcal{T}$ in the episode is obtained based on the definition in Equation (1, 2, 3). Then, store the tuple $\mathcal{T}$ in replay buffer. Each epoch can generate a kind of permutation of all documents $\{D^{k}\}_{k=1}^{N}$ in all training data. The second part is updating models with parameters $\Theta$ for $N\_update$ times. And the update is performed by sampling data through a mini-batch approach from $\mathcal{M}$. The whole network is trained end-to-end by minimizing the Temporal Difference (TD) (Sutton, 1988) loss shown in Equation (14) which is widely used in MARL to maximize $Q_{tot}$.
(15) $y^{tot}=\mathcal{R}+\gamma\max_{a^{\prime}}Q_{tot}(\textbf{o}^{\prime},\textbf{a}^{\prime},\bm{s}^{\prime};\Theta^{\prime})=\mathcal{R}$
@@ 5. Experiments
(RQ.2) How does MA4DIV perform on a larger scale industrial dataset compared to the baselines?
(RQ.3) How efficient is MA4DIV in training and inference process?
5.1. Datasets and Experimental Settings
We conduct experiments to address above research questions on two datasets, the public TREC 2009$\sim$2012 Web Track datasets and the industrial DU-DIV dataset. The details of these two datasets are shown in Table 1.
TREC 2009$\sim$2012 Web Track datasets are publicly available and Xia et al. (Xia et al., 2017) first uses them for training and evaluating diversified search models. Almost all the following works, such as $\mathrm{M^{2}Div}$ (Feng et al., 2018), DALETOR (Yan et al., 2021), MO4SRD (Yu, 2022), etc., conduct experiments on these datasets. As a widely used test collection for diversified search, the TREC Web track datasets only contains 198 queries, and previous studies often use Doc2vec (Le and Mikolov, 2014) to obtain the document embeddings on this dataset. In this work, to ensure a fair and meaningful comparison with existing work, we first compare MA4DIV with existing baselines on the this relatively small dataset. Then, we further conduct extensive experiments on a new and larger scale diversity dataset called DU-DIV, which is constructed based on real user data in industrial search engine. More details of datasets can be seen in Appendix D.
Table 1. Comparisons of datasets TREC and DU-DIV.
Dataset TREC DU-DIV
Source TREC 2009$\sim$2012 Web Track Real search engine
No. of Queries 198 4473
No. of Documents/Query average 211 docs/query 15 docs/query
No. of Subtopics maximum 7 subtopics/doc total 50 subtopics
Representation Doc2vec (Le and Mikolov, 2014) BERT (Devlin et al., 2018)
We conduct 5-fold cross-validation experiments on the reshaped TREC dataset with the same subset split based on unique queries as in (Feng et al., 2018). In the DU-DIV dataset, we randomly sample 80% of 4473 unique queries and their associated document lists to the training set, and the left 20% to the test set. The code framework of MA4DIV is based on PyMARL2 framework (Hu et al., 2021), which is a open-source code framework for multi-agent reinforcement learning. More detail settings of experiments are shown in Appendix E.
We compare MA4DIV with several state-of-the-art baselines in search result diversification, including:
5.2. Experimental Results
Table 2 and 3 report the performances of all baselines and our MA4DIV in terms of the six diversity performance metrics, including $\alpha$-NDCG@5, $\alpha$-NDCG@10, ERR-IA@5, ERR-IA@10, S-recall@5, S-recall@10 on the TREC web track datasets and the DU-DIV dataset respectively. The detail of evaluation metrics can be found in Appendix B. We highlight the top three algorithms on each metric by darkening the background color of the numbers.